OpenStudy (anonymous):

Rewrite 3x + 4y = 10 using function notation

OpenStudy (anonymous):

f(x)=10-3x/4

jimthompson5910 (jim_thompson5910):

Solve for y and then replace 'y' with f(x) \[\Large 3x + 4y = 10\] \[\Large 4y = 10-3x\] \[\Large 4y = -3x+10\] \[\Large y = \frac{-3x+10}{4}\] \[\Large y = \frac{-3x}{4}+\frac{10}{4}\] \[\Large y = -\frac{3}{4}x+\frac{5}{2}\] \[\Large f(x) = -\frac{3}{4}x+\frac{5}{2}\]